OpenStudy (anonymous):

-17, -12, -7, -2, ... Find an equation for the nth term of the arithmetic sequence.

OpenStudy (rea201):

Ok, so your writing an equation for the sequence right?

OpenStudy (anonymous):

yes

OpenStudy (rea201):

ok so what we do first is find the common difference

OpenStudy (rea201):

Do you know what the common difference is?

OpenStudy (anonymous):

5

OpenStudy (shiraz14):

Rea201 is correct ...

OpenStudy (rea201):

Correct so the function is going to be the first factor in this problem plus 5x. Do you know why it is 5x?

OpenStudy (anonymous):

You're multiplying it by something in the equation?

OpenStudy (rea201):

No. x represents the place where the nth term sits in the sequence. For example if I wanted to find the 10th term in the sequence that x=10. Because the term is in the 10th place

OpenStudy (anonymous):

okay

OpenStudy (rea201):

So do you know what the equation would look like?

OpenStudy (rea201):

@lornbeach

OpenStudy (anonymous):

an = -17 + 5(n - 1)

OpenStudy (rea201):

yes you are correct because a1=-17+5(1-1) a1=-17

OpenStudy (anonymous):

Yay thanks(:

OpenStudy (rea201):

No problem.